Suppose you borrowed $14, 000 at the rate of 10% and must repay in five equal annual installment at the end of each of the next five years. How much interest would you to have pay in the years 1-5?
1
Expert's answer
2019-12-20T03:47:57-0500

R = 14,000/5 = $2,800. — annual payment of principal.

Y1= (14,000 - 2,800 х 0) х 0,1 = $1,400

Y2 = (14,000 - 2,800 х 1) х 0,1 = $1,260

Y3= (14,000 - 2,800 х 2) х 0,1 = $1,148

Y4= (14,000 - 2,800 х 3) х 0,1 = $1,056

Y5 = (14,000 - 2,800 х 4) х 0,1 = $978